An object is placed between two plane mirrors inclined at an angle of 45° to each other. How many images do you expect to see?

Grade:10

8 Answers

sumit
7 years ago
NO OF IMAGES , WHEN THE OBJECT IS PLACED IN BETWEEN TWO PLANE MIRRORS IS GIVEN BY FORMULA =360/N-1
WHERE N IS ANGLE BETWEEN MIRRORS
360/45-1= 8-1=7 IMAGES 
unknown
7 years ago
 
NO OF IMAGES , WHEN THE OBJECT IS PLACED IN BETWEEN TWO PLANE MIRRORS IS GIVEN BY FORMULA =360/N-1
WHERE N IS ANGLE BETWEEN MIRRORS
360/45-1= 8-1=7 IMAGES 
Dr Vrinda Akshay
7 years ago
Number of images = 360/A - 1.
​in which A is the degree
360/45-1
8-1=7
​therefore the number of images is 7
Ritik Raj
7 years ago
I got the answer for this question.....pls solve this http://www.askiitians.com/forums/10-grade-science/when-secondary-and-tertiary-alcohol-is-oxidised-th_168199.htm
MD ADIL KARIM
7 years ago
360/angle given if the no of images is even we should subtract 1 from the no of images 
360/45=8 it is even  so ans is 7
Shahnawaz Khan
6 years ago
Using formula : 360/N - 1, Where N is degree in which two images are kept= 360/45 - 1 = 8 - 1 = 7Hence , seven images will be formed.
YASH
6 years ago
CcNo. Of images =(360/@)-1Here,@ stands for the angle. So, no. of images==8-1=7It implies that 7 images will be formed
praveen kumar
6 years ago
good question but it is easy when an object is placed between two plane mirrors inclined at an angle of 45° to each other the no of images is formed is given by the popular formula
no of images=(360/angle inclined)-1
there fore the answer for questions is 7 by this formula